OpenStudy (anonymous):

Show all steps used to solve the problem. 17-x/4=-10

OpenStudy (saifoo.khan):

-x/4 = -10 -17 -x/4 = -27 cancel the negatives of both sides, x/4 = 27 x = 27*4 x = Answer.

OpenStudy (anonymous):

x=108